# Holomorphic Yang-Mills fields on $B$-branes

### v1

@@ 3.3. Flag varieties
Let $T$ denote the maximal torus of ${\rm GL}(n,\,{\mathbb{C}})$ formed by the diagonal matrices. We set $W:=N(T)/T$, where $N(T)$ is the normalizer of $T$ in the group ${\rm GL}(n,\,{\mathbb{C}}).$ As it is well-known, the Weyl group $W$ is isomorphic to the symmetric group $S_{n}$. The Bruhat partial order [9, Sect 10.5] in $S_{n}$ will be denoted by $\leq$.
@@
We will set $C^{0}_{k}$ for the cell $C^{0}_{w_{k}}$ and $C_{k}:=C_{w_{k}}.$ The cell $C^{0}_{w}$ is an affine variety isomorphic to ${\mathbb{C}}^{l(\omega)}$ [9, page 157]. Hence, ${\rm dim}\,C_{4}=1={\rm dim}\,C_{5}$ and ${\rm dim}\,C_{2}=2={\rm dim}\,C_{3}.$ Furthermore, $X$ is the disjoint union of the cells
@@
$Z_{0}:=X\supset Z_{1}:=C_{2}\cup C_{3}\supset Z_{2}:=C_{4}\cup C_{5}\supset Z_{3}:=C_{6}.$
That is, $Z_{i}$ is the union of the Schubert varieties of dimension $3-i.$ We set $U_{i}:=Z_{i-1}\setminus Z_{i}$. From (3.16), it follows the expression of each $U_{i}$ in terms of Schubert cells. The corresponding inclusion relations are showed in the diagram.
[…]
@@
Let $Y$ be a complex variety, and $Y_{1},\,Y_{2}$ affine subvarieties of $Y$, such that $Y=Y_{1}\cup Y_{2}.$ If $Y_{k}$, for $k=1,2,$ is isomorphic to an affine variety, then the derived category $D^{b}(Y)$ is generated by
$\{\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,\alpha_{2!}{\mathscr{O}}_{Y_{2}}\}\cup\alpha_{!}\{{\rm generators\,of}\,D^{b}(Y_{1}\cap Y_{2})\},$
where $\alpha$ is the inclusion $\alpha:Y_{1}\cap Y_{2}\hookrightarrow Y,$ and $\alpha_{k}:Y_{k}\hookrightarrow Y$ denotes the corresponding embedding.
Proof. Given ${\mathscr{F}}\in\mathbf{Coh}(Y),$ one has the coherent sheaves ${\mathscr{F}}_{k}:=\alpha_{k!}\alpha_{k}^{!}{\mathscr{F}}$ and ${\mathscr{F}}_{Y_{1}\cap Y_{2}}=\alpha_{!}\alpha^{!}{\mathscr{F}}$ and the exact sequence of ${\mathscr{O}}_{Y}$-modules [22, Prop. 2.3.6 (vii)]
$0\to{\mathscr{F}}_{Y_{1}\cap Y_{2}}\to{\mathscr{F}}_{1}\oplus{\mathscr{F}}_{2}\to{\mathscr{F}}\to 0$
By Proposition 7, $D(Y_{k})$ is generated by ${\mathscr{O}}_{Y_{k}}.$ From the above exact sequence, it follows that the sheaf ${\mathscr{F}}$ belongs to the subcategory of $D^{b}(Y)$ generated by $\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,$ $\alpha_{2!}{\mathscr{O}}_{Y_{2}}$ together with the elements of form $\alpha_{!}(g),$ where $g$ belongs to a set of generators of $D^{b}(Y_{1}\cap Y_{2}).$
@@
Since $C^{0}_{4}\cap C^{0}_{5}=\emptyset,$ from Lemma 2 together with Corollary 2, it follows that the set $\{\alpha_{k!}{\mathscr{O}}_{C^{0}_{k}}\}_{k=4,5}$ is a set of generators of $D^{b}(U_{3}).$
For $k=4,5$ the ${\mathscr{O}}_{Z_{2}}$-module $(j_{3}\alpha_{k})_{!}{\mathscr{O}}_{C^{0}_{k}},$ will be denoted ${\mathscr{M}}_{k}.$
@@
For $r=1,\dots,6$
${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{r},\,{\mathscr{L}}_{r}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=0.$
Proof. As ${\mathscr{L}}_{1}={\mathscr{M}}_{1}$ is the invertible sheaf ${\mathscr{O}}_{X}(-Z_{1}),$
${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{1},\,{\mathscr{L}}_{1}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})={\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{O}}_{X},\,\Omega^{1}_{X})=H^{1,0}(X).$
By (3.17), $H^{1,0}(X)=0.$
The remaining cases can be treated similarly. To simplify the notation we will write ${\mathscr{X}}$ for ${\mathscr{O}}_{X}$ and ${\mathscr{Z}}_{1}$ for ${\mathscr{O}}_{Z_{1}};$ we also write $i:=i_{1}.$
As $\Omega^{1}_{X}$ is a locally free ${\mathscr{X}}$-module, by [22, Prop. 2.5.13]
${\mathscr{L}}_{2}\otimes_{\mathscr{X}}\Omega^{1}_{X}\simeq i_{!}\big{(}{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}.$
Hence,
${\mathscr{H}om}_{\mathscr{X}}\big{(}{\mathscr{L}}_{2},\,{\mathscr{L}}_{2}\otimes_{\mathscr{X}}\Omega^{1}_{X}\big{)}\simeq{\mathscr{H}om}_{\mathscr{X}}\big{(}i_{*}{\mathscr{M}}_{2},\,i_{!}({\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}})\big{)}$
$\simeq i_{*}{\mathscr{H}om}_{{\mathscr{Z}}_{1}}\big{(}{\mathscr{M}}_{2},\,{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}.$
Since ${\mathscr{M}}_{2}$ is an invertible ${\mathscr{Z}}_{1}$-module
${\rm Hom}_{\mathscr{X}}\big{(}{\mathscr{L}}_{2},\,{\mathscr{L}}_{2}\otimes_{\mathscr{X}}\Omega^{1}_{X}\big{)}\simeq\Gamma\big{(}Z_{1},\,{\mathscr{H}om}_{{\mathscr{Z}}_{1}}\big{(}{\mathscr{M}}_{2},\,{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}\big{)}$
$={\rm Hom}_{{\mathscr{Z}}_{1}}\big{(}{\mathscr{M}}_{2},\,{\mathscr{M}}_{2}\otimes_{{\mathscr{Z}}_{1}}\Omega^{1}_{Z_{1}}\big{)}=H^{0}(Z_{1},\,\Omega^{1}_{Z_{1}})$
On the other hand, $Z_{1}=C_{2}\cup C_{3}$, and $H^{1,0}(C_{2})=H^{1,0}(C_{3})=0.$ Thus, $H^{0}(Z_{1},\,\Omega^{1}_{Z_{1}})=0.$ That is, the proposition is proved for the case $r=2.$ The proof of the remaining cases is analogous. ∎
@@
Proof. An object of $D^{b}(X)$ is isomorphic to one of the triangulated category generated by the family ${\mathscr{L}}_{1},\dots,{\mathscr{L}}_{6}$ in $D^{b}(X).$ Let $({G}^{\bullet},\,d^{\bullet})$ be an object of this triangulated category. Then ${G}^{p}$ is a sheaf of the form
(3.22) ${G}^{p}=\bigoplus_{i\in S_{p}}{\mathscr{L}}_{pi},$
$S_{p}$ being a finite set. From Corollary 1 together with (3.21) and Proposition 10, it follows
${\rm Hom}_{D^{b}(X)}\big{(}{G}^{\bullet},\,\Omega^{1}(G^{\bullet})\big{)}=0.$
The theorem is a consequence of Proposition 5. ∎
3.4. The homotopy category
Let $X$ be a smooth projective variety. The category $\mathbf{Coh}(X)$ of coherent sheaves over $X$ has not enough injectives, for this reason it is convenient to regard $D^{b}(X)$ as a subcategory of $D^{b}(\mathbf{Mod}({\mathscr{O}_{X}}))$, the bounded derived category of the ${\mathscr{O}}_{X}$-modules. In fact, $D^{b}(X)$ is equivalent to $D^{b}(\mathbf{Mod}({\mathscr{O}_{X}}))_{\rm coh}$, the full subcategory of $D^{b}(\mathbf{Mod}({\mathscr{O}_{X}}))$ consisting of the complexes with coherent cohomology [29, Exp II] [15, Chapter II] (for a brief description [19, page 63]). Thus, $D^{b}(X)$ can be identified with the homotopy category whose objects are the complexes ${\mathscr{G}}^{\bullet}$ of injective ${\mathscr{O}}_{X}$-modules, such that its cohomogy is bounded and coherent; i.e. ${\mathscr{H}}^{j}({\mathscr{G}}^{\bullet})$ is coherent and vanishes for $|j|>>0$. We will denote this homotopy category by $K^{b}(X)_{\rm coh}$.
In the identification of $D^{b}(X)$ with the homotopy category $K^{b}(X)_{\rm coh},$ the quasi-isomorphisms correspond to homotopy equivalences. Thus, the independence of quasi-isomorphisms in the frame of $K(X)_{\rm coh}$ means independence with respect to homotopy equivalences. In particular, if ${\mathscr{F}}^{\bullet}$ and ${\mathscr{G}}^{\bullet}$ are objects of $K^{b}(X)_{\rm coh}$ homotopically equivalent, the respective cohomology sheaves ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})$ and ${\mathscr{H}}^{i}({\mathscr{G}}^{\bullet})$ are isomorphic.
If $Y$ is a Hodge manifold, then it is a smooth projective variety, according to a well-known Kodaira’s theorem. By the GAGA correspondence, the coherent analytic sheaves on $Y$ can be identified with the algebraic ones. Hence, the $B$-branes on $Y$ can be considered as objects of $K^{b}(Y)_{\rm coh}$.
From now on, we assume that $Y$ is a Hodge manifold. In accordance with the preceding paragraph, one can assume that the $B$-brane $({\mathscr{F}}^{\bullet},\delta^{\bullet})$ over $Y$ is a complex of injective ${\mathscr{O}}$-modules with coherent cohomology modules satisfying ${\mathscr{H}}^{i}({\mathscr{F}}^{\bullet})=0$ for $|i|>>0$. Moreover,
${\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\simeq{\rm Hom}_{K^{b}(Y)_{\rm coh}}({\mathscr{F}}^{\bullet},\,\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}})),$
where $\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}})$ is an object of $K^{b}(Y)_{\rm coh}$ $q$-isomorphic to ${\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})$. Thus, the elements of the space (3.1) can be identified with morphisms in the homotopyy category $K^{b}(Y)_{\rm coh}$. Given a gauge $\psi\in{\rm Hom}_{D^{b}(Y)}({\mathscr{F}}^{\bullet},\,{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})),$ we denote by $\hat{\psi}$ the corresponding morphism in the category $K^{b}(Y)_{\rm coh}.$ The exact sequence (3.4) gives rise to the following sequence of arrows in the additive category $K^{b}(Y)_{\rm coh}$
$\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}\overset{\hat{\iota}}{\rightarrow}\widehat{{\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})}\overset{\hat{\pi}}{\rightarrow}{\mathscr{F}}^{\bullet}$
where $\widehat{\Omega^{1}({\mathscr{F}}^{\bullet})}$ is quasi-isomorphic to $\Omega^{1}({\mathscr{F}}^{\bullet})$ and $\hat{\pi}\circ\hat{\psi}=1.$ We will take to the cohomology of those complexes taking into account that:
  • The cohomology of an object with a hat is isomorphic to the cohomology of the corresponding complex without hat.
  • A morphism of the homotopy category, determines a well-defined morphism between the cohomologies; i. e. a morphism of complexes of ${\mathscr{O}}$-modules.
  • ${\mathscr{H}}^{*}(\Omega^{1}({\mathscr{F}}^{\bullet}))\simeq\Omega^{1}({\mathscr{H}}^{*}({\mathscr{F}}^{\bullet})).$
Then, one has
(3.23) $0\to\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))\overset{\boldsymbol{\iota}^{j}}{\longrightarrow}{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\overset{\boldsymbol{\pi}^{j}}{\longrightarrow}{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to 0,$
where the maps induced between the cohomologies are denoted in bold. In turn, $\hat{\psi}$ determines a well-defined morphism of ${\mathscr{O}}$-modules between the cohomologies
$\boldsymbol{\psi}^{j}:{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\to{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})),$
that obviously satisfies $\boldsymbol{\pi}^{j}\circ\boldsymbol{\psi}^{j}=1.$
Hence, the sequence (3.23) of ${\mathscr{O}}$-modules splits. Furthermore, by (3.5), ${\mathscr{J}}^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))\simeq{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet})).$ Thus, we have a splitting, in the category of ${\mathscr{O}}$-modules, of ${\mathscr{J}}^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))$ as direct sum ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})$ and $\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})).$
According to Definition 1, $\boldsymbol{\psi}^{j}$ defines a holomorphic connection on the ${\mathscr{O}}$-module ${\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}).$
We set $\boldsymbol{\eta}^{j}$ for the morphism of abelian sheaves defined by the inclusion in the direct sum
$\boldsymbol{\eta}^{j}:\sf{Forg}\big{(}{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\big{)}\rightarrow\sf{Fog}\big{(}{\mathscr{H}}^{j}({\mathscr{J}}^{1}({\mathscr{F}}^{\bullet}))\big{)}=\sf{Forg}\big{(}{\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})\oplus\Omega^{1}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet}))\big{)},$
Hence, $\boldsymbol{\pi}^{j}(\boldsymbol{\psi}^{j}-\boldsymbol{\eta}^{j})=0$, and thus $\boldsymbol{\psi}^{j}-\boldsymbol{\eta}^{j}$ defines a morphism of abelian sheaves
(3.24) $\vartheta^{j}:{\mathscr{H}}^{j}({\mathscr{F}})\to{\Omega^{1}}({\mathscr{H}}^{j}({\mathscr{F}}^{\bullet})),$
which, by (3.2), satisfies the Leibniz’s rule (2.14). That is,

### v2

@@ 2.3.2. Flag varieties
Let $T$ denote the maximal torus of ${\rm GL}(n,\,{\mathbb{C}})$ formed by the diagonal matrices. We set $W:=N(T)/T$, where $N(T)$ is the normalizer of $T$ in the group ${\rm GL}(n,\,{\mathbb{C}}).$ As it is well-known, the Weyl group $W$ is isomorphic to the symmetric group $S_{n}$. The Bruhat partial order [8, Sect. 10.5] in $S_{n}$ will be denoted by $\leq$.
@@
We will set $C^{0}_{k}$ for the cell $C^{0}_{w_{k}}$ and $C_{k}:=C_{w_{k}}.$ The cell $C^{0}_{w}$ is an affine variety isomorphic to ${\mathbb{C}}^{l(\omega)}$ [8, p. 157]. Hence, ${\rm dim}\,C_{4}=1={\rm dim}\,C_{5}$ and ${\rm dim}\,C_{2}=2={\rm dim}\,C_{3}.$ Furthermore, $X$ is the disjoint union of the cells
@@
$X=Z_{0}=C_{1}\overset{i_{1}}{\hookleftarrow}Z_{1}=C_{2}\cup C_{3}\overset{i_{2}}{\hookleftarrow}Z_{2}=C_{4}\cup C_{5}\overset{i_{3}}{\hookleftarrow}Z_{3}=C_{6}$
That is, $Z_{i}$ is the union of the Schubert varieties of dimension $3-i.$ We set $U_{i}:=Z_{i-1}\setminus Z_{i}$. From (2.17), it follows the expression of each $U_{i}$ in terms of Schubert cells. One has the the following inclusion relations
$Z_{0}\overset{j_{1}}{\hookleftarrow}U_{1}=C^{0}_{1},\;\;\;Z_{1}\overset{j_{2}}{\hookleftarrow}U_{2}=C^{0}_{2}\cup C^{0}_{3},\;\;\;Z_{2}\overset{j_{3}}{\hookleftarrow}U_{3}=C^{0}_{4}\cup C^{0}_{5}.$
@@
Let $Y$ be a complex variety, and $Y_{1},\,Y_{2}$ disjoint affine subvarieties of $Y,$ then the derived category $D^{b}(Y)$ is generated by $\{\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,\alpha_{2!}{\mathscr{O}}_{Y_{2}}\},$ where $\alpha_{k}:Y_{k}\hookrightarrow Y$ denotes the corresponding embedding.
Proof. Given ${\mathscr{F}}\in\mathbf{Coh}(Y),$ one has the coherent sheaves ${\mathscr{F}}_{k}:=\alpha_{k!}\alpha_{k}^{!}{\mathscr{F}}$ and the exact sequence of ${\mathscr{O}}_{Y}$-modules [16, Prop. 2.3.6 (vii)]
$0\to{\mathscr{F}}_{Y_{1}\cap Y_{2}}=0\to{\mathscr{F}}_{1}\oplus{\mathscr{F}}_{2}\to{\mathscr{F}}\to 0$
By Proposition 3, $D(Y_{k})$ is generated by ${\mathscr{O}}_{Y_{k}}.$ From the above exact sequence, it follows that the sheaf ${\mathscr{F}}$ belongs to the subcategory of $D^{b}(Y)$ generated by $\alpha_{1!}{\mathscr{O}}_{Y_{1}},\,$ $\alpha_{2!}{\mathscr{O}}_{Y_{2}}.$
@@
The ${\mathscr{O}}_{Z_{2}}$-module $(j_{3}\alpha_{k})_{!}{\mathscr{O}}_{C^{0}_{k}},$ will be denoted ${\mathscr{M}}_{k}.$
@@
${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{r},\,{\mathscr{L}}_{r}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=0,\,$ for $r=1,\dots,6.$
Proof. As in the proof of Lemma 2, ${\rm Hom}_{{\mathscr{O}}_{X}}({\mathscr{L}}_{1},\,{\mathscr{L}}_{1}\otimes_{{\mathscr{O}}_{X}}\Omega^{1}_{X})=H^{1,0}(Z_{1}).$ On the other hand, $Z_{1}=C_{2}\cup C_{3}$, and by (2.18) $H^{1,0}(C_{2})=H^{1,0}(C_{3})=0.$ Thus, $H^{0}(Z_{1},\,\Omega^{1}_{Z_{1}})=0.$ That is, the proposition is proved for the case $r=1.$ The proof of the remaining cases is analogous. ∎
@@
Proof. An object of $D^{b}(X)$ is isomorphic to one of the triangulated category generated by the family ${\mathscr{L}}_{1},\dots,{\mathscr{L}}_{6}$ in $D^{b}(X).$ Let $({G}^{\bullet},\,d^{\bullet})$ be an object of this triangulated category. Then ${G}^{p}$ is a sheaf of the form $G^{p}=\oplus_{i}{\mathscr{L}}_{pi},$ with $pi\in\{1,\dots,6\}.$ From Corollary 1 together with (2.22) and Proposition 6, it follows ${\rm Hom}_{D^{b}(X)}\big{(}{G}^{\bullet},\,\Omega^{1}(G^{\bullet})\big{)}=0.$ The theorem is a consequence of Proposition 1. ∎
3. Holomorphic Yang-Mills fields
3.1. Hermitian sheaves
The fiber at $x\in X$ of a coherent ${\mathscr{O}}_{X}$-module ${\mathscr{G}}$ will be denoted by ${\mathscr{G}}_{(x)}:={\mathscr{G}}_{x}/{\mathfrak{m}}_{x}{\mathscr{G}}_{x}$, where ${\mathfrak{m}}_{x}$ is the maximal ideal of ${\mathscr{O}}_{x}$. If ${\mathcal{Z}}$ is a section of ${\mathscr{G}}$ the corresponding vector in ${\mathscr{G}}_{(x)}$ is denoted by $Z(x)$. On the other hand, the singular set ${\mathcal{S}}$ of ${\mathscr{G}}$ is a closed analytic subset of $X$ whose codimension is greater than or equal to $1$ [19, Chap. V, Thm. 5.8]. Moreover, ${\mathscr{G}}$ is locally free on $X\setminus{\mathcal{S}}$. We set ${G}$ for the vector bundle over $X\setminus{\mathcal{S}}$, with fibers $G(x):={\mathscr{G}}_{(x)}$, determined by the locally free sheaf ${\mathscr{G}}|_{X\setminus{\mathcal{S}}}$.
The following definition generalizes the given in [29, Chap. III, Sect. 1] for locally free sheaves.
Definition 3.
A Hermitian metric on the coherent sheaf ${\mathscr{G}}$ is a set $\{\langle\,,\,\rangle_{x}\}_{x\in X}$ of Hermitian metrics on the fibers of ${\mathscr{G}}$, satisfying the following condition: Given ${\mathcal{Z}}_{1},{\mathcal{Z}}_{2}$ two sections of ${\mathscr{G}}$ on an open $U$ of $X,$ the map
Given ${\mathcal{Z}}_{1},{\mathcal{Z}}_{2}\in\Gamma(X,\,{\mathscr{G}}),\,$ ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ is a bounded map with possible discontinuities in the singular locus. If $X$ is a Kähler manifold, the function ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ may be non-integrable on $X$ with respect to the volume form defined by the Kähler metric.
If ${\mathscr{G}}$ is a torsion-free coherent sheaf on the Kähler manifold $X,$ then ${\rm codim}\,{\mathcal{S}}\geq 2$ [19, Cap. V, Cor. 5.15]. Hence, there exists an analytic closed subset ${\mathcal{W}}$ of $X,$ containing ${\mathcal{S}},$ and such that $1\leq{\rm codim}\,{\mathcal{W}}<{\rm codim}\,{\mathcal{S}}.$ Let ${\mathcal{N}}$ be a tubular neighborhood of ${\mathcal{S}}$ in ${\mathcal{W}}.$ Then
  1. (1)
    ${\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}$ is a continuous map on the closed space $X\setminus{\mathcal{N}},$ which is disjoint with the singularity set ${\mathcal{S}}.$
  2. (2)
    ${\mathcal{W}},$ as a subset of $X$ with codimension $\geq 1,$ has measure zero with respect to the differential form ${\rm d}{\rm vol}.$
Thus, the following integral is well-defined
(3.1) $({\mathcal{Z}}_{1},\,{\mathcal{Z}}_{2}):=\int_{X\setminus{\mathcal{N}}}{\mathcal{Z}}_{1}\cdot{\mathcal{Z}}_{2}\,{\rm d}{\rm vol}.$
Remark 5.
[…]
3.1.1. Yang-Mills functional
Given $\nabla$ a holomorphic connection on the coherent sheaf $\mathscr{F}.$ Then the corresponding operator $\nabla$ defines a morphism of ${\mathbb{C}}_{X}$-modules $\nabla^{(k)}:\Omega^{k}({\mathscr{F}}){\to}\Omega^{k+1}({\mathscr{F}})$ in the usual way. The composition ${\mathcal{K}}_{\nabla}:=\nabla^{(1)}\circ\nabla:{\mathscr{F}}\to\Omega^{2}({\mathscr{F}})$ is the curvature of $\nabla$; and as it is well-known
(3.2) ${\mathcal{K}}_{\nabla}\in{\rm Hom}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))=\Gamma(X,\,{\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))).$
The connection is said to be flat if ${\mathcal{K}}_{\nabla}=0$.
For each $x\in X,$ we denote by $\alpha_{x}$ and $\lambda_{x}$ the natural morphisms
$\big{(}{\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))\big{)}_{x}\overset{\alpha_{x}}{\to}{\rm Hom}_{{\mathscr{O}}_{x}}\big{(}{\mathscr{F}}_{x},\,\Omega^{k}_{x}\otimes_{{\mathscr{O}}_{x}}{\mathscr{F}}_{x}\big{)}\overset{\lambda_{x}}{\leftarrow}\Omega^{k}_{x}\otimes_{{\mathscr{O}}_{x}}{\rm End}_{{\mathscr{O}}_{x}}({\mathscr{F}}_{x}).$
As ${\mathscr{F}}$ is coherent, $\alpha_{x}$ is isomorphism [10, p. 239]. Furthermore, if ${\mathscr{F}}_{x}$ is free, then $\lambda_{x}$ is bijective.
If ${\mathcal{S}}$ is the singularity set of ${\mathscr{F}},$ it is also the singular locus of the sheaf ${\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}})).$ Hence, for each point $x$ outside of ${\mathcal{S}},$ the fiber of ${\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))$ at $x$ can be identified with the vector space $\Omega^{k}{(x)}\otimes{\rm End}({F}{(x)}).$ According to (3.2), the curvature ${\mathcal{K}}_{\nabla}$ of a holomorphic connection $\nabla$ determines the vector
(3.3) $K_{\nabla}(x)\in\Omega^{2}(x)\otimes{\rm End}(F(x))$
for each $x\in X\setminus{\mathcal{S}}$. That is, $K_{\nabla}$ is a $2$-form ${\rm End}(F)$-valued.
Let us assume that ${\mathscr{F}}$ is a Hermitian torsion-free sheaf on the Kähler manifold $X,$ according to (3.1), one defines
(3.4) $\lVert{\mathcal{K}}_{\nabla}\rVert^{2}=({\mathcal{K}}_{\nabla},\,{\mathcal{K}}_{\nabla})=\int_{X\setminus{\mathcal{N}}}{\mathcal{K}}_{\nabla}\cdot{\mathcal{K}}_{\nabla}\,{\rm d}{\rm vol}=\int_{X\setminus{\mathcal{N}}}|K_{\nabla}\wedge\star\,K_{\nabla}|,$
[…]
More concretely, if locally $K_{\nabla}$ can be expressed as $\alpha\otimes A$, with $\alpha$ a $2$-form and $A$ a local section of ${\rm End}(F)$, then the integrand in (3.4) is $(\alpha\wedge\star\alpha)\langle A\circ A\rangle$. In a local unitary frame of ${\rm End}(F)$, if the connection is compatilble with the metric, the matrix $\check{A}$ associated to $A$ is antihermitian and $\langle A\circ A\rangle=(-1/2){\rm tr}(\check{A}\check{A})$. That is,
(3.5) $|K_{\nabla}\wedge\star\,K_{\nabla}|=-(1/2){\rm tr}\big{(}K_{\nabla}\wedge\star\,K_{\nabla}\big{)}.$
On the space of holomorphic gauge fields on the torsion-free sheaf ${\mathscr{F}}$ one can define the following map $\,\mathcal{YM}:\nabla\in{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))\mapsto\lVert{\mathcal{K}}_{\nabla}\rVert^{2}.\,$ It is called the Yang-Mills’ functional. The $\nabla$ on which this functional takes a stationary value are the holomorphic Yang-Mills fields.
3.1.2. Yang-Mills fields on reflexive sheaves
If ${\mathscr{G}}$ is a reflexive sheaf on $X,$ then it is torsion-free and the codimension of the singular set is $\geq 3$ [19, Cap. V, Cor. 5.20], [13, Cor. 1.4]. Furthermore, if $C$ is a closed subset of $X$ with codimension $\geq 2,$ then the restriction $\Gamma(X,\,{\mathscr{G}})\to\Gamma(X\setminus C,\,{\mathscr{G}})$ is an isomorphism [19, Cap. V, Prop. 5.21] [14, Prop. 1.11].
Let us suppose that ${\mathscr{F}}$ is a reflexive sheaf on the Kähler manifold $X$ endowed with a Hermitian metric. Then ${\mathscr{E}nd}({\mathscr{F}})$ is also a reflexive sheaf [19, Chap. V, Prop. 4.15]. If ${\mathcal{S}}$ is the singular locus of ${\mathscr{F}},$ let ${\mathcal{W}}$ be a closed subspace of $X$ containing ${\mathcal{S}}$ and such that ${\rm codim}\,{\mathcal{W}}=2.$ Let ${\mathcal{N}}$ denote a tubular neighborhood of ${\mathcal{S}}$ in ${\mathcal{W}}.$ Therefore, the vector space $\Gamma(X\setminus{\mathcal{N}},\,\Omega^{k}({\mathscr{E}nd}({\mathscr{F}})))\simeq\Gamma(X,\,\Omega^{k}({\mathscr{E}nd}({\mathscr{F}}))),$ and the latter is finite dimensional. Furthermore, as $(X\setminus{\mathcal{N}})\cap{\mathcal{S}}=\emptyset,\,$ ${\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\,\Omega^{k}({\mathscr{F}}))|_{X\setminus{\mathcal{N}}}$ is a locally free sheaf.
Let $\nabla$ be a holomorphic gauge field on ${\mathscr{F}}$, then
${\mathcal{K}}_{\nabla}\in\Gamma\big{(}X,\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))\big{)}=\Gamma(X\setminus{\mathcal{N}},\,{\mathscr{H}om}({\mathscr{F}},\,\Omega^{2}({\mathscr{F}}))\big{)}$
can considered as a global section of the locally free sheaf $\big{(}\Omega^{2}\otimes{\rm End}(F)\big{)}|_{X\setminus{\mathcal{S}}}.$ A similar observation is also valid for $\nabla$.∇ .
Let us assume that the sheaf ${\mathscr{F}}$ supports a holomorphic connection $\nabla_{0}$. By Proposition 1, given ${\mathcal{E}}_{1},\dots\mathcal{E}_{m}$, a basis of
(3.6) ${\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}}))=\Gamma\big{(}Y,\,{\mathscr{H}om}_{{\mathscr{O}}_{X}}({\mathscr{F}},\Omega^{1}({\mathscr{F}}))\big{)}=\Gamma(Y\setminus{\mathcal{S}},\,\Omega^{1}\otimes{\rm End}(F)),$
where ${\mathcal{B}}_{i}:=\nabla_{0}({\mathcal{E}}_{i})$ and ${\mathcal{B}}_{ij}:={\mathcal{E}}_{i}\wedge{\mathcal{E}}_{j}$. If moreover ${\mathscr{F}}$ hermitian, then
where $P$ is a polynomial of degree $\leq 4$ in the variables $\lambda_{i}$.
(3.8) $\frac{\partial\,P}{\partial\lambda_{i}}=0,\;\;\;i=1,\dots,m.$
Therefore,
Theorem 3.
If the Hermitian reflexive sheaf ${\mathscr{F}}$ admits a holomorphic gauge field and $m={\rm dim}\,{\rm Hom}({\mathscr{F}},\,\Omega^{1}({\mathscr{F}})).$ Then the set ${\sf YM}({\mathscr{F}})$ of holomorphic Yang-Mills fields on ${\mathscr{F}}$ is in bijective correspondence with the points of the algebraic set in ${\mathbb{C}}^{m}$ defined by $m$ algebraic equations.
Proof. The case $m=2$ is a consequence of Bézout’s theorem. ∎
Taking into account Proposition 1 and (3.6), any “variation” of a holomorphic connection $\nabla$ on ${\mathscr{F}}$ can be written as $\nabla_{\epsilon}=\nabla+\epsilon E,$ with $\epsilon\in{\mathbb{C}}$ and $E\in\Gamma(X\setminus S,\,\Omega^{1}\otimes{\rm End}(F)).$
(3.9) $(1/2)\frac{d}{d\epsilon}\Big{|}_{\epsilon=0}||K_{\nabla_{\epsilon}}||^{2}=\int_{X\setminus{\mathcal{N}}}\langle K_{\nabla},\,\nabla E\rangle\,{\rm dvol}=:(K_{\nabla},\,\nabla{E}).$
(3.10) $(K_{\nabla},\,\nabla{E})=0.$
In particular, the flat holomorphic gauge fields are Yang-Mills.
We denote
${}^{(p)}\nabla:\Gamma(X,\,\Omega^{p}({\mathscr{E}nd}({\mathscr{F}})))\to\Gamma(X,\,\Omega^{p+1}({\mathscr{E}nd}({\mathscr{F}}))),$
${}^{(2)}\nabla K_{\nabla}=0.$